The ratio of boys to girls in a math class is 2 : 3. There are 18 girls in
the class. How many students are in the class?

Answer by stanbon(75887)

The ratio of boys to girls in a math class is 2 : 3. There are 18 girls in
the class. How many students are in the class?
------------
Let 2x be number of boys.
Then 3x is number of girls.
---
Solve: 3x = 18
x = 6
---
# of students = 2x+3x = 5x
5x = 5*6 = 30 students/
============================
Cheers,
Stan H.